# ODGEN: Domain-specific Object Detection Data Generation with Diffusion Models

### v1

@@ 1 Introduction
Object detection is one of the most widely used computer vision tasks in real-world applications. However, data scarcity poses a significant challenge to the performance of state-of-the-art models like YOLOv5 [23] and YOLOv7 [52]. With the progress of generative diffusion models, such as Stable Diffision [42], and DALL-E [12], which can generate high-quality images from text prompts, recent works have started to explore the potential of synthesizing images for perceptual model training. Furthermore, methods like ControlNet [60], ReCo [59], and GLIGEN [31] provide various ways to control the contents of synthetic images. Concretely, extra visual or textual conditions are introduced as guidance for diffusion models to generate objects under certain spatial constraints. Therefore, it becomes feasible to generate images together with instance-level or pixel-level annotations.
Nevertheless, it remains challenging to generate an effective supplementary training set for real-world object detection applications. Firstly, large-scale pre-trained diffusion models are usually trained on web crawl datasets such as LAION [46], whose distributions may be quite distinct from specialist domains. The domain gap could undermine the fidelity of generated images, especially the detailed textures and styles of the objects, e.g., the tumor in medical images or the avatar in a video game. Secondly, the text prompts for object detection scenes may contain multiple categories of subjects. It could cause the "concept bleeding" [38] problem for modern diffusion models, which is defined as the unintended merging or overlapping of distinct visual elements in an image. As a result, the synthetic images can be inconsistent with pseudo labels used as generative conditions. Thirdly, overlapping objects, which are common in object detection datasets, are likely to be merged as one single object by existing controllable generative methods. Lastly, some objects may be neglected by the diffusion model so that no foreground is generated at the conditioned location. All of the limitations can hinder the downstream object detection performance.
@@ 2 Related Works
Diffusion models [47, 19, 48, 7, 36, 27] have made great progress in recent years and become mainstream for image generation, showing higher fidelity and training stability than prior generative models like GANs [15, 4, 25, 26, 24] and VAEs [28, 41, 50]. Latent diffusion models (LDMs) [42] perform the denoising process in the compressed latent space and achieve flexible generators conditioned on inputs like text and semantic maps. LDMs significantly improve computational efficiency and enable training on internet-scale datasets [46]. Modern large-scale text-to-image diffusion models including Imagen [44], DALL·E [40], eDiff-I [2], and Stable Diffusion [42, 38] have demonstrated unparalleled capabilities to produce diverse samples with high fidelity given unfettered text prompts.
@@
Dataset synthesis for training object detection models. Copy-paste [9] is a simple but effective data augmentation method for detectors. Simple Copy-Paste [13] achieves improvements with a simple random placement strategy for objects. Following works [12, 63] generate foreground objects and then paste them on background images. However, generated images by these methods usually have obvious artifacts on the boundary of pasted regions. DatasetGAN [62] takes an early step to generate labeled images automatically. Diffusion models have been used to synthesize training data and benefit downstream tasks including object detection [11, 5, 61, 10], image classification [1, 8, 17, 32, 45, 49], and semantic segmentation [54, 21, 14, 29, 37, 55, 58, 34, 57]. Modern approaches for image-label pairs generation can be roughly divided into two categories. One group of works [14, 54, 11, 29, 34, 57, 61] first generate images and then apply a pre-trained perception model to generate pseudo labels on the synthetic images. The other group [5, 37, 21] uses the same strategy as our approach to synthesize images under the guidance of annotations as input conditions. The second group also overlaps with layout-to-image approaches [31, 60, 59, 65, 53, 22, 64, 30]. Our approach should be assigned to the second group and is designed to address challenging cases like multi-class objects, occlusions between objects, and specific domains.
@@ 3.2 Object-wise Conditioning
ControlNet [60] can perform controllable synthesis with visual conditions. However, the control can be challenging with an increasing object number and category number due to the "concept bleeding" phenomenon [38]. Stronger conditions are needed to ensure high-fidelity generation. To this end, we propose a novel object-wise conditioning strategy with ControlNet.
Text list encoding. As shown in Fig. 2 (b) and (c), given a set of object class names and their bounding boxes, we first build a text list consisting of each object with the fixed template "a <classname>". Then the text list is padded with empty texts to a fixed length $N$ and converted to a list of embeddings by a pre-trained CLIP text tokenizer and encoder. The embeddings are stacked and encoded by a 4-layer convolutional text embedding encoder, and used as the textual condition for ControlNet. The text encoding of native ControlNet compresses the information in global text prompts altogether, resulting in mutual interference between distinct categories. To alleviate this "concept bleeding" [38] problem of multiple categories, our two-step text encoding enables the ControlNet to capture the information of each object with separate encoding.
@@ 3.3 Dataset Synthesis Pipeline for Object Detection
4 Experiments
We conduct extensive experiments to demonstrate the effectiveness of the proposed ODGEN in both specific and general domains. FID [18] is used as the metric of fidelity. Mean average precisions (mAP) of YOLOv5s [23] and YOLOv7 [52] trained with synthetic data are used to evaluate the trainability, which concerns the usefulness of synthetic data for detector training. Our approach is implemented with Stable Diffusion v2.1 and compared with prior controllable generation methods based on Stable Diffusion, including ReCo [59], GLIGEN [31], ControlNet [60], GeoDiffusion [5], and MIGC [65]. Native ControlNet doesn’t support bounding box conditions. Therefore, we convert boxes to a mask $C$ sized $H\times W\times K$, where $H$ and $W$ are the image height and width, and $K$ is the class number. If one pixel $(i,j)$ is in $n$ boxes of class $k$, the $C[i,j,k]=n$. YOLO models are trained with the same recipe as Roboflow-100 [6] for 100 epochs to ensure convergence. The length of text and image lists $N$ is set to the maximum object number per image in the training set.
@@ 4.1 Specific Domains
Roboflow-100 [6] is used for the evaluation in specific domains. It consists of 100 object detection datasets of various domains. We select 7 representative datasets (RF7) covering video games, medical imaging, and underwater scenes. To simulate data scarcity, we only sample 200 images as the training set for each dataset. We fine-tune all methods with the 200 images and generate 5000 new synthetic images for evaluation. $\lambda$ in Eq. 1 and $\gamma$ in Eq. 2 are set as 1 and 25. We first fine-tune the diffusion model according to Fig. 2 (a) for 3k iterations. Then we train the object-wise conditioning modules on a V100 GPU with batch size 4 for 200 epochs, the same as the other methods to be compared.
@@ 4.1.1 Fidelity
For each dataset in RF7, we compute the FID scores of synthetic images against real images, respectively. As shown in Tab. 1, our approach outperforms all the other methods. We visualize the generation quality in Fig. 4. GLIGEN only updates gated self-attention layers inserted to Stable Diffusion, making it hard to be generalized to specific domains like MRI images and the Apex game. ReCo and GeoDiffusion integrate encoded bounding box information into text tokens, which require more data for diffusion model and text encoder fine-tuning. With only 200 images, they fail to generate objects in the given box regions. ControlNet, integrating the bounding box information into the visual condition, presents more consistent results with the annotation. However, it may still miss some objects or generate wrong categories in complex scenes. ODGEN achieves superior performance on both visual effects and consistency with annotations.
@@ 4.1.2 Trainability
4.2 General Domains
COCO-2014 [35] is used to evaluate our method in general domains. As the scenes in this dataset are almost covered by the pre-training data of Stable Diffusion, we skip the domain-specific fine-tuning stage and train the diffusion model together with the object-wise conditioning modules. We train our ODGEN on the COCO training set with batch size 32 for 60 epochs on ×$\times 8$ V100 GPUs as well as GLIGEN, ControlNet, and GeoDiffusion. Officially released checkpoints for ReCo (trained for 100 epochs) and MIGC (trained for 300 epochs) are directly employed for comparison. $\gamma$ in Eq. 2 is set as 10 for our ODGEN trained on COCO.
@@ 4.2 General Domains
Metrics ReCo GLIGEN ControlNet GeoDiffusion MIGC ODGEN
FID 18.36 26.15 25.54 30.00 21.82 16.16
mAP@.50 7.60 / 11.01 6.70 / 9.42 1.43 / 1.15 5.94 / 9.21 9.54 / 16.01 18.90 / 24.40
mAP@.50:.95 3.82 / 5.29 3.56 / 4.60 0.52 / 0.38 2.37 / 4.44 4.67 / 8.65 9.70 / 14.20
@@ 4.2.1 Fidelity
We use the annotations of the COCO validation set (41k) as conditions to generate 41k images. The FID scores in Tab. 3 are computed against real images in the COCO validation set. ODGEN achieves better FID results than the other methods. We provide a typical visualized example in Fig. 5. Given overlapping bounding boxes of multiple categories, ODGEN synthesizes all the objects of correct categories and accurate positions, outperforming the other methods. Substantial visualized samples are added in Appendix F.
@@ Appendix C Supplemental Ablations
Detectors training w/o real data. In the main paper, we train the detectors with 200 real images and 5000 synthetic images for RF7 benchmarks. In this ablation study, we conduct experiments with the 5000 synthetic images only, to explore the impact of the absence of real data for training detectors. As shown in Tab. 7, for easy datasets with fewer objects, like MRI image and Cotton, training with 5000 synthetic data only achieves better results than training with 200 real samples only. Because the generation quality is good enough in these cases, increasing the dataset size can improve the model performance. Nevertheless, for complex scenes like Robomaster and Aquarium, training with synthetic data only fails to achieve better results. It is caused by the limited generation quality of modern diffusion models. For instance, the VAE in Stable Diffusion is not good at synthesizing texts in images, which are common in the Robomaster dataset. Besides, FID results in Tab. 1 indicate that there still exist gaps between the distributions of real and synthetic images. The performance of our approach is promising to be further improved in the future with more powerful generative diffusion models.
@@
Table 8: Ablations of category isolation during datasets synthesis (Metric: mAP@.50 of YOLOv5s / YOLOv7, real + synth images # 200 + 5000). It hinders the performance in most cases.
Category Isolation True False
@@
Underwater 41.1 / 42.0 40.0 / 44.8

### v2

@@ 1 Introduction
Object detection is one of the most widely used computer vision tasks in real-world applications. However, data scarcity poses a significant challenge to the performance of state-of-the-art models like YOLOv5 [25] and YOLOv7 [54]. With the progress of generative diffusion models, such as DALL-E [13], and Stable Diffision [44], which can generate high-quality images from text prompts, recent works have started to explore the potential of synthesizing images for perceptual model training. Furthermore, methods like GLIGEN [33], ReCo [61], and ControlNet [62] provide various ways to control the contents of synthetic images. Concretely, extra visual or textual conditions are introduced as guidance for diffusion models to generate objects under certain spatial constraints. Therefore, it becomes feasible to generate images together with instance-level or pixel-level annotations.
Nevertheless, it remains challenging to generate an effective supplementary training set for real-world object detection applications. Firstly, large-scale pre-trained diffusion models are usually trained on web crawl datasets such as LAION [48], whose distributions may be quite distinct from specialist domains. The domain gap could undermine the fidelity of generated images, especially the detailed textures and styles of the objects, e.g., the tumor in medical images or the avatar in a video game. Secondly, the text prompts for object detection scenes may contain multiple categories of subjects. It could cause the "concept bleeding" [12, 40, 68] problem for modern diffusion models, which is defined as the unintended merging or overlapping of distinct visual elements in an image. As a result, the synthetic images can be inconsistent with pseudo labels used as generative conditions. Thirdly, overlapping objects, which are common in object detection datasets, are likely to be merged as one single object by existing controllable generative methods. Lastly, some objects may be neglected by the diffusion model so that no foreground is generated at the conditioned location. All of the limitations can hinder the downstream object detection performance.
@@ 2 Related Works
Diffusion models [7, 20, 29, 38, 49, 50] have made great progress in recent years and become mainstream for image generation, showing higher fidelity and training stability than prior generative models like GANs [4, 16, 26, 27, 28] and VAEs [30, 43, 52]. Latent diffusion models (LDMs) [44] perform the denoising process in the compressed latent space and achieve flexible generators conditioned on inputs like text and semantic maps. LDMs significantly improve computational efficiency and enable training on internet-scale datasets [48]. Modern large-scale text-to-image diffusion models including eDiff-I [2], DALL·E [42], Imagen [46], and Stable Diffusion [40, 44] have demonstrated unparalleled capabilities to produce diverse samples with high fidelity given unfettered text prompts.
@@
Dataset synthesis for training object detection models. Copy-paste [9] is a simple but effective data augmentation method for detectors. Simple Copy-Paste [14] achieves improvements with a simple random placement strategy for objects. Following works [13, 65] generate foreground objects and then paste them on background images. However, generated images by these methods usually have obvious artifacts on the boundary of pasted regions. DatasetGAN [64] takes an early step to generate labeled images automatically. Diffusion models have been used to synthesize training data and benefit downstream tasks including object detection [5, 10, 11, 63], image classification [1, 8, 18, 34, 47, 51], and semantic segmentation [15, 23, 31, 36, 39, 56, 57, 59, 60]. Modern approaches for image-label pairs generation can be roughly divided into two categories. One group of works [11, 15, 31, 36, 56, 59, 63] first generate images and then apply a pre-trained perception model to generate pseudo labels on the synthetic images. The other group [5, 23, 39] uses the same strategy as our approach to synthesize images under the guidance of annotations as input conditions. The second group also overlaps with layout-to-image approaches [24, 33, 55, 61, 62, 66, 67]. Our approach should be assigned to the second group and is designed to address challenging cases like multi-class objects, occlusions between objects, and specific domains.
@@ 3.2 Object-wise Conditioning
ControlNet [62] can perform controllable synthesis with visual conditions. However, the control can be challenging with an increasing object number and category number due to the "concept bleeding" phenomenon. Stronger conditions are needed to ensure high-fidelity generation. To this end, we propose a novel object-wise conditioning strategy with ControlNet.
Text list encoding. As shown in Fig. 2 (b) and (c), given a set of object class names and their bounding boxes, we first build a text list consisting of each object with the fixed template "a <classname>". Then the text list is padded with empty texts to a fixed length $N$ and converted to a list of embeddings by a pre-trained CLIP text tokenizer and encoder. The embeddings are stacked and encoded by a 4-layer convolutional text embedding encoder, and used as the textual condition for ControlNet. The text encoding of native ControlNet compresses the information in global text prompts altogether, resulting in mutual interference between distinct categories. To alleviate this "concept bleeding" problem of multiple categories, our two-step text encoding enables the ControlNet to capture the information of each object with separate encoding.
@@ 3.3 Dataset Synthesis Pipeline for Object Detection
4 Experiments
We conduct extensive experiments to demonstrate the effectiveness of the proposed ODGEN in both specific and general domains. FID [19] is used as the metric of fidelity. Mean average precisions (mAP) of YOLOv5s [25] and YOLOv7 [54] trained with synthetic data are used to evaluate the trainability, which concerns the usefulness of synthetic data for detector training. Our approach is implemented with Stable Diffusion v2.1 and compared with prior controllable generation methods based on Stable Diffusion, including ReCo [61], GLIGEN [33], ControlNet [62], GeoDiffusion [5], InstanceDiffusion [55], and MIGC [67]. Native ControlNet doesn’t support bounding box conditions. Therefore, we convert boxes to a mask $C$ sized $H\times W\times K$, where $H$ and $W$ are the image height and width, and $K$ is the class number. If one pixel $(i,j)$ is in $n$ boxes of class $k$, the $C[i,j,k]=n$. YOLO models are trained with the same recipe as Roboflow-100 [6] for 100 epochs to ensure convergence. The length of text and image lists $N$ used in our ODGEN is set to the maximum object number per image in the training set.
@@ 4.1 Specific Domains
Roboflow-100 [6] is used for the evaluation in specific domains. It consists of 100 object detection datasets of various domains. We select 7 representative datasets (RF7) covering video games, medical imaging, and underwater scenes. To simulate data scarcity, we only sample 200 images as the training set for each dataset. The whole training process including the fine-tuning on both cropped objects and entire images and the training of the object-wise conditioning module, only depends on the 200 images. $\lambda$ in Eq. 1 and $\gamma$ in Eq. 2 are set as 1 and 25. We first fine-tune the diffusion model according to Fig. 2 (a) for 3k iterations. Then we train the object-wise conditioning modules on a V100 GPU with batch size 4 for 200 epochs, the same as the other methods to be compared.
@@ 4.1.1 Fidelity
For each dataset in RF7, we compute the FID scores of 5000 synthetic images against real images, respectively. As shown in Tab. 1, our approach outperforms all the other methods. We visualize the generation quality in Fig. 4. GLIGEN only updates gated self-attention layers inserted to Stable Diffusion, making it hard to be generalized to specific domains like MRI images and the Apex game. ReCo and GeoDiffusion integrate encoded bounding box information into text tokens, which require more data for diffusion model and text encoder fine-tuning. With only 200 images, they fail to generate objects in the given box regions. ControlNet, integrating the bounding box information into the visual condition, presents more consistent results with the annotation. However, it may still miss some objects or generate wrong categories in complex scenes. ODGEN achieves superior performance on both visual effects and consistency with annotations.
@@ 4.1.2 Trainability
In addition, we add experiments with larger-scale datasets with 1000 images sampled from the Apex Game and the Underwater datasets. The training setups are kept the same as the training on 200 images. We conduct experiments on ODGEN, ReCo, and GeoDiffusion. ReCo and GeoDiffusion may benefit from larger-scale training datasets since they need to fine-tune more parameters in both the UNet in Stable Diffusion and the CLIP text encoder. GLIGEN struggles to adapt to new domains. ControlNet performs worse on layout control than ODGEN. Therefore, these two methods are not included in this part. The corrupted label filtering step is not used for any method. Results in Tab. 3 show that the baselines trained on real data only become stronger with larger-scale datasets while our ODGEN still benefits detectors with synthetic data and outperforms other methods.
Table 3: mAP@.50 / mAP@.50:.95 ($\uparrow$) results of ODGEN trained on larger-scale datasets of 1000 real images. The top 3 rows show results of YOLOv5s and the bottom 3 rows show results of YOLOv7. Baseline models are trained with 1000 real images only, whereas the other models are trained with 1000 real + 5000 / 10000 synthetic images from various methods. ODGEN leads to more significant improvement than other methods.
Datasets Apex Game Apex Game Apex Game Underwater Underwater Underwater
real + synth # 1000 + 0 1000 + 5000 1000 + 10000 1000 + 0 1000 + 5000 1000 + 10000
ReCo 83.2 / 53.5 78.7 / 46.9 82.0 / 46.9 55.6 / 29.2 55.1 / 28.4 55.9 / 29.1
GeoDiffusion 83.2 / 53.5 80.0 / 47.2 82.5 / 47.5 55.6 / 29.2 54.2 / 27.9 54.3 / 28.0
ODGEN 83.2 / 53.5 83.3 / 53.5 83.6 / 53.6 55.6 / 29.2 59.6 / 32.5 56.3 / 29.8
ReCo 83.8 / 55.0 80.5 / 50.7 79.2 / 49.9 54.6 / 28.3 56.5 / 28.7 56.4 / 30.1
GeoDiffusion 83.8 / 55.0 81.2 / 51.0 81.0 / 50.5 54.6 / 28.3 57.0 / 28.9 55.8 / 28.9
ODGEN 83.8 / 55.0 84.4 / 55.2 84.0 / 55.0 54.6 / 28.3 58.2 / 29.8 62.1 / 31.8
@@
Metrics ReCo GLIGEN ControlNet Geo- Diffusion MIGC Instance- Diffusion ODGEN
FID 18.36 26.15 25.54 30.00 21.82 23.29 16.16
mAP@.50 7.60 / 11.01 6.70 / 9.42 1.43 / 1.15 5.94 / 9.21 9.54 / 16.01 10.00 / 17.10 18.90 / 24.40
mAP@.50:.95 3.82 / 5.29 3.56 / 4.60 0.52 / 0.38 2.37 / 4.44 4.67 / 8.65 5.42 / 10.20 9.70 / 14.20
4.2 General Domains
COCO-2014 [37] is used to evaluate our method in general domains. As the scenes in this dataset are almost covered by the pre-training data of Stable Diffusion, we skip the domain-specific fine-tuning stage and train the diffusion model together with the object-wise conditioning modules. We train our ODGEN on the COCO training set with batch size 32 for 60 epochs on ×$\times 8$ V100 GPUs as well as GLIGEN, ControlNet, and GeoDiffusion. Officially released checkpoints of ReCo (trained for 100 epochs on COCO), MIGC (trained for 300 epochs on COCO), and InstanceDiffusion (trained on self-constructed datasets) are employed for comparison. $\gamma$ in Eq. 2 is set as 10 for our ODGEN.
Table 5: mAP@.50 ($\uparrow$) and mAP@.50:.95 ($\uparrow$) of YOLOv5s / YOLOv7 on the COCO dataset. Baseline models are trained with 80k images from the COCO training set, whereas the other models are trained with the same 80k real + 20k synthetic images. ODGEN outperforms baseline and the other methods.
Metrics Baseline ReCo GLIGEN ControlNet
mAP@.50 51.5 / 64.5 50.8 / 64.3 50.9 / 64.2 50.1 / 64.3
mAP@.50:.95 32.6 / 45.4 31.8 / 45.2 31.9 / 45.2 31.0 / 45.2
Metrics GeoDiffusion MIGC InstanceDiffusion ODGEN
mAP@.50 51.2 / 64.4 51.5 / 64.6 51.5 / 64.6 52.1 / 65.0
mAP@.50:.95 32.1 / 45.2 32.5 / 45.5 32.6 / 45.6 33.1 / 45.9
@@ 4.2.1 Fidelity
We use the annotations of the COCO validation set as conditions to generate 41k images. The FID scores in Tab. 4 are computed against the COCO validation set. ODGEN achieves better FID results than the other methods. We provide a typical visualized example in Fig. 5. Given overlapping bounding boxes of multiple categories, ODGEN synthesizes all the objects of correct categories and accurate positions, outperforming the other methods. More samples are added in Appendix G.
@@ 4.2.2 Trainability
We further conduct experiments by adding 20k synthetic images to the 80k training images. We train YOLO models from scratch on the COCO training set (80k images) as the baseline and on the same 80k real images + 20k synthetic images generated by different methods for comparison. We use the labels of 20k images from the COCO validation set as conditions to generate the synthetic set and use the other 21k real images for evaluation. The results are shown in Tab. 5. It shows that ODGEN improves the mAP@.50:95 by 0.5% and outperforms the other methods.
We observe that YOLO models trained on synthetic data only fall behind models trained on real data. We add experiments with different training and validation data combinations in Tab. 6. YOLO models trained on real images show better generalization ability and achieve close results when tested on real and synthetic data. YOLO models trained on synthetic data only get significantly better results when tested on synthetic data than when generalized to real data. It indicates that noticeable domain gaps still exist between real and synthetic data, which may be limited by the generation quality of modern Stable Diffusion models and are promising to be narrowed with future models.
Table 6: mAP@.50 ($\uparrow$) and mAP@.50:.95 ($\uparrow$) of YOLOv5s / YOLOv7 trained from scratch and validated on real or synthetic COCO validation set. 10k images are used for training and the other 31k images are used for validation. Real represents real images in the COCO validation set and synthetic represents images synthesized by our ODGEN using the same labels.
Train Validate mAP@.50 ($\uparrow$) mAP@.50:95 ($\uparrow$) Train Validate mAP@.50 ($\uparrow$) mAP@.50:95 ($\uparrow$)
Real Real 29.40 / 41.70 15.80 / 26.30 Synthetic Real 18.90 / 24.40 9.70 / 14.20
Real Synthetic 29.40 / 39.90 15.00 / 23.20 Synthetic Synthetic 37.90 / 45.40 20.40 / 27.10
@@ Appendix C Supplemental Ablations
Detector training w/o real data. In the main paper, we train the detectors with 200 real images and 5000 synthetic images for RF7 benchmarks. In this ablation study, we conduct experiments with the 5000 synthetic images only to explore the impact of the absence of real data for training detectors. As shown in Tab. 10, for easy datasets with fewer objects, like MRI image and Cotton, training with 5000 synthetic data only achieves better results than training with 200 real samples only. Increasing the dataset size can improve the model performance because the generation quality is good enough in these cases. Nevertheless, for complex scenes like Robomaster and Aquarium, training with synthetic data only fails to achieve better results. It may be caused by the limited generation quality of modern diffusion models. For instance, the VAE in Stable Diffusion is not good at synthesizing texts in images, which is common in the Robomaster dataset. Besides, FID results in Tab. 1 indicate that there still exist gaps between the distributions of real and synthetic images. The performance of our approach is promising to be further improved in the future with more powerful generative diffusion models. In addition, the current dataset synthesis pipeline is not fully optimized and cannot reproduce the distributions of foreground objects perfectly, which may also lead to worse performance of training on synthetic data only with the RF7 datasets.
@@
Number of synthesized samples. We ablate the number of synthesized samples used for detector training and provide quantitative results in Tab. 12. With increasing amounts of synthesized samples, detectors achieve higher performance and get very close results between 5000 and 10000 synthetic samples.
@@
Table 11: Ablations of the number of synthetic samples (Metric: mAP@.50 ($\uparrow$) of YOLOv5s / YOLOv7, real images # 200). Results are very close between 5k and 10k synthetic samples.
synth # Robomaster Cotton Aquarium
0 50.3 / 48.8 29.4 / 32.0 53.8 / 53.0
1000 52.2 / 61.7 44.0 / 51.3 55.0 / 63.8
3000 61.9 / 61.1 62.6 / 56.4 56.8 / 65.8
5000 67.4 / 65.4 63.9 / 55.0 62.6 / 66.5
10000 67.4 / 65.2 64.4 / 56.5 62.8 / 66.3
Table 12: Ablations of category isolation during datasets synthesis (Metric: mAP@.50 ($\uparrow$) of YOLOv5s / YOLOv7, real + synth images # 200 + 5000). It hinders the performance in most cases.
Split Categories True False
@@
Underwater Object 41.1 / 42.0 40.0 / 44.8
Corrupted label filtering. We design this step to filter some labels when objects are not generated successfully, which may be caused by some unreasonable boxes obtained with the pipeline in Fig. 3. For experiments on COCO, this step is not applied to any method since we directly use labels from the COCO validation set and hope to synthesize images consistent with the real-world labels. For RF7 experiments, this step is only applied to our ODGEN in Tab. 2. For fair comparison, we skip this step on RF7 to compare the generation capability of different methods fairly and provide results in Tab. 13. It shows that the corrupted label filtering step only contributes a small part of the improvement. Without this step, our method still outperforms the other methods significantly. In addition, as illustrated in Appendix A, the current dataset synthesis pipeline is designed to compare different methods and can be improved further in future work.
Table 13: mAP@.50:.95 ($\uparrow$) of YOLOv5s / YOLOv7 on RF7. Baseline models are trained with 200 real images only, whereas the other models are trained with 200 real + 5000 synthetic images from various methods. The corrupted label filtering step is not applied to all methods. ODGEN still leads to the biggest improvement on all 7 domain-specific datasets.
Baseline ReCo GLIGEN ControlNet GeoDiffusion ODGEN
real + synth # 200 + 0 200 + 5000 200 + 5000 200 + 5000 200 + 5000 200 + 5000
Apex Game 38.3 / 47.2 25.0 / 31.5 24.8 / 32.5 33.8 / 42.7 29.2 / 35.8 39.8 / 52.6
Robomaster 27.2 / 26.5 18.2 / 27.9 19.1 / 25.0 24.4 / 32.9 18.2 / 22.6 39.0 / 33.3
Cotton 16.7 / 20.5 29.3/ 37.5 28.0 / 39.0 22.6 / 35.1 30.2 / 36.0 40.5 / 42.1
Road Traffic 35.3 / 41.0 22.8 / 29.3 22.2 / 29.5 22.1 / 30.5 17.2 / 29.4 38.2 / 43.2
Aquarium 30.0 / 29.6 23.8 / 34.3 24.1 / 32.2 18.2 / 25.6 21.6 / 30.9 32.0 / 38.4
Underwater 16.7 / 19.4 13.7 / 15.8 14.9 / 18.5 15.5 / 17.8 13.8 / 17.2 18.9 / 21.6
Foreground region enhancement on COCO. We set $\gamma$ as 10 in Eq. 2 in our experiments trained on COCO. We add ablations with $\gamma$ value of 0 to ablate foreground region enhancement on COCO and provide quantitative results in Tab. 14. It shows that the foreground region enhancement helps ODGEN achieve better results, especially the layout-image consistency reflected by mAP results.
Table 14: FID ($\downarrow$) and mAP ($\uparrow$) of YOLOv5s / YOLOv7 on COCO. FID is computed with 41k synthetic images. For mAP, YOLO models are trained from scratch on 10k synthetic images and validated on 31k real images. Foreground region enhancement contributes to the improvement of both FID and mAP results.
Method FID ($\downarrow$) mAP@.50 ($\uparrow$) mAP@.50:95 ($\uparrow$)
ODGEN w/ foreground region enhancement 16.16 18.90 / 24.40 9.70 / 14.20
ODGEN w/o foreground region enhancement 16.45 16.90 / 23.10 8.72 / 13.50
Fine-tuning on both cropped foreground objects and entire images. For specific domains, taking the Robomaster dataset in RF7 as an example, most images in the dataset contain multiple categories of objects like "armor", "base", "watcher", and "car", which are either strange for Stable Diffusion or different from what Stable Diffusion tends to generate with the same text prompts. For models fine-tuned on entire images only, they cannot obtain direct guidance on which parts in entire images correspond to these objects. As a result, the fine-tuned models cannot synthesize correct images for foreground objects given text prompts like "a base in a screen shot of the Robomaster game". We provide FID results of foreground objects synthesized by models fine-tuned on both entire images and cropped foreground objects (text prompts are composed of object names and the scene name, e.g., "a car and a base in a screen shot of the Robomaster game") and models fine-tuned on entire images only in Tab. 15. It shows that the proposed approach helps fine-tuned models generate images of foreground objects significantly better. Similarly, we also provide FID results of the Road Traffic dataset in Tab. 16, which is relatively more familiar for Stable Diffusion than the Robomaster dataset. Results also show the improvement provided by our approach compared with models fine-tuned on entire images only. Visualized examples produced by the model fine-tuned on entire images only are shown in Fig. 8. Such models struggle to capture the information of foreground objects and fail to produce images needed to build image lists for our ODGEN.
Table 15: FID ($\downarrow$) results of foreground objects in the Robomaster dataset synthesized by models fine-tuned on different data.
Object categories watcher armor car base rune
Entire images only 384.94 532.56 364.46 325.79 340.11
Entire images and cropped objects 124.45 136.27 135.66 146.50 128.95
Table 16: FID ($\downarrow$) results of foreground objects in the Road Traffic dataset synthesized by models fine-tuned on different data.
Object categories vehicle traffic light motor- cycle fire hydrant cross- walk bus bicycle
Entire images only 241.64 240.55 213.90 205.22 253.40 238.63 153.76
Entire images and cropped objects 105.27 90.71 143.30 53.05 97.33 118.58 65.61
Figure 8: Visualized examples produced by the model fine-tuned on entire images from the Road Traffic dataset only. It fails to produce images of foreground objects correctly.